OpenStudy (anonymous):

The value of Peter's house has risen 30% a year for the last 3 years. If the value of Peter's home 3 years ago was $130,000, what is the value now?

OpenStudy (razzanide):

130,000 X 0.3 = 39,000, 130,000 + 39,000 =169,000 for the value after the first year 169,000 X 0.3 = 50,700, 169,000 + 50,700 =219,700 for the value after the second year 219,700 X 0.3 = 65,910, 219,700 + 65,910 =285,610 for the value after the third year I think this is right?